# Addressing the gravitational collapse of a massless scalar field with Physics-Informed Neural Networks

### v1

@@ 4 Results
Training curves. As noted earlier, the relative quadratic errors $l_{2}$ defined in Eq. (38) can be decomposed into a sum of components, $l_{2}=l_{2}^{\phi}+l_{2}^{\alpha}+l_{2}^{C}$. This decomposition allows us to assess the performance of each variable individually, rather than relying solely on the total sum, which may lead to masked deficiencies in training. In Figure 3, we present the learning curves for these metrics across all considered problems, focusing on the four best performing models. For each variable, the error scale is kept fixed in all different scenarios. The first conclusion is that the cases showing gravitational collapse, especially the critical case ($\phi_{0}=0.1125$), are by far the most challenging to train, achieving higher errors in general and evolving within a much narrower range. In general, all models achieve comparable performance in this case. It is worth highlighting that, given the construction of the $l_{2}$ metric, small variations can translate into substantial improvements in the target variable—particularly in regions exhibiting sharp changes over small $(t,r)$ intervals. Moreover, even though most models converge faster than the vanilla PINN, none stands out as decisively superior—except ModPINN, which yields an average total error of 0.588, compared with 0.694 and 0.698 for the next best (the vanilla and PirateNet, respectively) and with variability roughly three times smaller. In principle, given the difficulty of the critical case, likely due to a sharp and hard-to-navigate loss landscape, we might expect independent trainings to yield slightly different results. However, as Table 1 has shown, the variability is concerning across models, on the order of $\sim 30\%$ relative to the mean, except for ModPINN, where it is only $\sim 10\%$. This indicates a more stable model that may admit further optimization and potentially even better results with a more exhaustive preconditioning of the hyperparameters. For the supercritical case, $\phi_{0}=0.125$, it seems that all models identify black hole formation more clearly as the dominant result, as evidenced by the errors being uniformly lower than those of the vanilla PINN. In addition, the learning curves suggest that the problem becomes less difficult with a substantially wider range of progress during training. Finally, for the subcritical cases, all models give similar solutions, except for $\alpha(t,r)$ produced by the vanilla PINN at $\phi_{0}=0.10625$. We also observe that every alternative model converges faster—roughly in half the training time—representing a clear advantage. However, in general, the main conclusion is that all methods perform reasonably well due to the relative “ease” of these problems: smooth gradients make them straightforward for any of the models to solve.
@@
It is important to note that these learning curves correspond to the training runs reported in table 1; thus, all methodological settings are the same, including the spatiotemporal sampling size $N_{t}\times N_{r}=128\times 128$. Several factors could be adjusted to improve results: increasing sampling density, fine-tuning $\epsilon_{t}$ and $\epsilon_{r}$, and /or refining the weights assigned to the initial and boundary conditions. For the critical case in particular, a more detailed analysis is needed, which we will present in the remainder of the paper.
@@
Robustness of ModPINN. Up to this point, we have reviewed a set of recent models in the literature, but it is worth putting more focus on our own proposal, which we have denoted as ModPINN. In essence, this architecture is no more than a modified PINN that replaces linear layers with quadratic ones, and whose main strength is an embedding introduced before the layer concatenation. As we have shown, this embedding leads to improved results for the supercritical case and, specifically, for the critical case $\phi_{0}=0.1125$, improving average performance and reducing variability. Taking into account that this model is the one giving the strongest results for gravitational collapse, it is necessary to carry on a more thorough examination of how various pre-training design choices influence its performance. Among these factors, we include the sampling size (that is, the number of spatiotemporal sampling points, $N_{t}\times N_{r}$) and the scale of the neural architecture, in particular the number of quadratic layers and the number of neurons per layer. Another parameter of interest that we examine here is the strength of the temporal causality, $\epsilon_{t}$, while $\epsilon_{r}$ will be constant and fixed at 10 for this analysis.
@@
  • The $10\times 75$ architecture—the deepest and widest configuration—yields the most stable overall results across all cases, much like the largest sampling size, $256\times 256$, which also exhibits the greatest stability. Consequently, the model clearly benefits from larger sampling grids and higher network capacities, providing robustness independent of the problem considered. In contrast, smaller networks display greater variability as $\epsilon_{t}$ changes. This behavior is a clear indication of the inherent physical complexity of the problem.
  • In general, the motivation behind the enforcement of temporal causality is to prevent PINNs from “looking ahead” in time, that is, from resolving late-time points before early-time ones [Wang_2024]. This approach can be helpful in problems with a strong dependence on initial conditions, where the evolution of the system is fully determined by them. Consequently, it is reasonable to expect that in a smooth, easy-to-solve case like $\phi_{0}=0.1$, causality enforcement will have limited impact, which is essentially what we observe. For this problem, there is no clear trend in $l_{2}$ as $\epsilon_{t}$ increases, because the solutions with $\epsilon_{t}=0$ are generally good and any observed improvements are not substantial.
  • Following the same logic, for the critical case $\phi_{0}=0.1125$, we observe a stronger dependence on $\epsilon_{t}$, although the errors are generally higher due to the steep highly localized gradients mentioned. We see that the model clearly benefits from enforcing causality, particularly for the $128\times 128$ case, where the error quickly drops for $\epsilon_{t}=10^{2}$. Regarding the $\phi_{0}=0.125$ scenario—specifically with $256\times 256$ points and a $10\times 75$ network—the solution is markedly stable, showing only minor sensitivity to $\epsilon_{t}$. By contrast, when smaller networks are used, the improvement becomes appreciable.
@@
Figure 4: Relative total $l_{2}$ error of the proposed ModPINN for the considered scalar-field initial-value problems (by column) as a function of the hyperparameter $\epsilon_{t}$, which modulates the strength of the temporal causality. Each row corresponds to a different spatiotemporal grid resolution $(t,r)$, and each sub-figure reports results for three network sizes.
@@
In the two bottom panels, we depict how the $l^{2}$ errors vary with the hyperparameter $\Lambda$, which controls the “strength” with which spacetime points $(t,r)$ are moved towards regions exhibiting higher values of the normalized physical loss, as defined in (28). The lower-left panel reports this magnitude for the two subcritical cases ($\phi_{0}=0.1,0.10625$), while the lower-right panel presents the corresponding results for the critical and supercritical cases ($\phi_{0}=0.1125,0.125$). We observe that as $\Lambda$ increases within the range considered, the error shows a clear decreasing trend for the subcritical scenarios, with a particular reduction in variability at $\Lambda=10$ for the case $\phi_{0}=0.1$. For the critical scenario $\phi_{0}=0.1125$, the trend is also downward from $\Lambda=0$ to 1, with the error dropping from approximately $\sim 1.0$ to $\sim 0.6$. This constitutes a substantial improvement that should not be eclipsed by the use of a broader scale relative to the other panel. Regarding $\Lambda=10$ we observe a sudden increase in error for this case, indicating that the solutions deteriorate when this parameter is overly enlarged. This behavior is somewhat expected: the physics lies on the boundary between criticality, and the collocation points are required and used within a highly localized region of spacetime mainly determined by the black hole formation. Excessive and rapid relocation of the points can move some of them into undesirable regions, thereby devaluating their utility and, overall, leading to worse solutions. Finally, for the supercritical case with $\phi_{0}=0.125$, the trend is essentially flat. It is worth noting that the model under consideration here is ModPINN, which—consistent with table 1—provides the best performance for this regime with quite low variability. Consequently, this scenario is already well resolved without additional techniques, as the model alone effectively minimizes the loss. It is therefore unsurprising that the solution does not improve substantially when adaptive sampling is introduced. By contrast, performing the same analysis with models such as the vanilla PINN—where this case is less accurately captured—might reveal a decreasing trend under the said sampling methodology.
@@ 5 Conclusions
The work reported in this paper has revisited this challenging physical problem, focusing on solutions at and around the critical point. Our primary aim has been to demonstrate that neural network-based methodologies, such as Physics-Informed Neural Networks (PINNs), can serve as a viable and useful alternative to finite differences. To this end, we have considered four different initial-value problems for the field, using the amplitude of the initial kernel, $\phi_{0}$, as the control parameter, with the critical case included among them. We have addressed these cases using several recent methodologies from the literature and introduced our own, termed ModPINN, which has been tailored to this class of scenarios, including several novel aspects from recent works such as coordinate embedding, quadratic neural layers, and learnable activation functions. Our model delivers slightly better critical solutions with better variability than other state-of-the-art approaches, even though all of them have led to very competitive predictions. Furthermore, by implementing a relatively simple algorithm that moves the spacetime sampling points $(t,r)$, we treat them as particles subject to a “force” with a prescribed acceleration modulated via the $\Lambda$ hyperparameter in Eq. (28), which indicates how quickly points migrate to regions with higher physical loss. The key consequence is that we are able to work with comparatively small sets, unlike traditional numerical methods that resolve black holes by brute force—highly densifying the mesh in spacetime regions that are known a priori to be most challenging.
@@
Irrespective of the specific models used, much of the substantive effort resides in the design of the physical loss. The boundary conditions for $r\gg 0$ introduced in Eq. (16), are formulated in terms of the derivatives of the physical variables, whose enforcement leads to the asymptotic behaviors that these variables are expected to have far from the source of perturbation. These conditions have proved to be crucial, as well as the mechanisms of temporal and spatial enforcements, which in some sense help to emulate the adaptive refinements of finite-difference schemes without incurring additional computational effort. In the analysis displayed in Figure 4, we have examined the robustness of ModPINN with respect to several factors, including the density of sampling points and the width and depth of the network. We have found that wider and deeper architectures indeed facilitate better solutions, and they also perform relatively good on comparatively large spatiotemporal grids such as $256\times 256$ points—still small in comparison with those used in finite-difference schemes. Overall, these configurations exhibit greater robustness and smoother performance across independent trials. While further analysis could be considered, this allows us to highlight a notable advantage of PINNs: aside from hardware constraints, there is substantial flexibility to scale model size and sampling, and the resulting $N_{t}\times N_{r}$ grids in deep learning setups are orders a lot smaller than those typically required by finite differences.
@@
Across all models considered for the target problems, we can conclude that physics-informed deep-learning approaches are indeed capable of resolving gravitational collapse driven by a massless scalar field in a spherically symmetric, time-dependent spacetime. By judiciously selecting activation functions to restrict variables to their physically feasible ranges and by formulating boundary conditions in derivative form, we obtain solutions close to the ground truth. While there is still room for improvement, the results indicate that these models are progressing in the right direction. Nevertheless, a PINN can be seen as an alternative (or surrogate) to a finite-difference numerical solver. Broadly speaking, PINN-type models employ the governing PDEs as components of a constructed loss function—alongside additional information—to solve the physical problem at hand. Consequently, once the initial scalar field is specified, the full evolution of $\phi$, $\alpha$, and $C$ over $(t,r)$ is determined. Thus the problem becomes deterministic; gravitational collapse may or may not occur depending on that initial configuration. Therefore, this imposes certain limitations when investigating aspects such as the universality of the exponent $\gamma$ in the scaling law of the black hole mass $M_{\mathrm{BH}}$ near criticality, as well as the generality of the proportionality constant $c_{f}$ with respect to the family of the initial field. To examine these aspects—and to study the self-similarity of solutions as $p\to p^{*}$—we would need access to a broad set of initial condition sets and their full corresponding solutions. Therefore, addressing these questions will require a different point of view—one that allows us to traverse families of initial conditions with full flexibility and to generate data with relative ease. In this sense, the present project paves the way to future studies that can probe the aforementioned issues, extend to alternative spacetime metrics and matter models, and tackle more complex scenarios in which finite-difference methods could be either inapplicable or impractical due to prohibitive computational costs.
@@ 6 Acknowledgments
This work has been partially funded by the Valencian Government grant with Reference Number CIAICO/2024/111; the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call-QuantumSpain project, and the European Union through the Recovery, Transformation and Resilience Plan—NextGenerationEU within the framework of the Digital Spain 2025 Agenda. This work has been also supported by the Spanish Agencia Estatal de Investigación (grant PID2024-159689NB-C21) funded by MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe, by the Prometeo excellence programme grant CIPROM/2022/49 funded by the Generalitat Valenciana, and by the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA2021-SE-01 Grant No. NewFunFiCO-101086251. MWC is supported by NSERC.

### v2

@@ 4 Results
Training curves. As noted earlier, the relative quadratic errors $l_{2}$ defined in (38) can be decomposed into a sum of components, $l_{2}=l_{2}^{\phi}+l_{2}^{\alpha}+l_{2}^{C}$. This decomposition allows us to assess the performance of each variable individually, rather than relying solely on the total sum, which may lead to masked deficiencies in training. In figure 3, we present the learning curves for these metrics across all considered problems, focusing on the four best performing models. For each variable, the error scale is kept fixed in all different scenarios. The first conclusion is that the cases showing gravitational collapse, especially the critical case ($\phi_{0}=0.1125$), are by far the most challenging to train, achieving higher errors in general and evolving within a much narrower range. In general, all models achieve comparable performance in this case. It is worth highlighting that, given the construction of the $l_{2}$ metric, small variations can translate into substantial improvements in the target variable—particularly in regions exhibiting sharp changes over small $(t,r)$ intervals. Moreover, even though most models converge faster than the vanilla PINN, none stands out as decisively superior—except ModPINN, which yields an average total error of 0.588, compared with 0.694 and 0.698 for the next best (the vanilla and PirateNet, respectively) and with variability roughly three times smaller. In principle, given the difficulty of the critical case, likely due to a sharp and hard-to-navigate loss landscape, we might expect independent trainings to yield slightly different results. However, as Table 1 has shown, the variability is concerning across models, on the order of $\sim 30\%$ relative to the mean, except for ModPINN, where it is only $\sim 10\%$. This indicates a more stable model that may admit further optimization and potentially even better results with a more exhaustive preconditioning of the hyperparameters. For the supercritical case, $\phi_{0}=0.125$, it seems that all models identify black hole formation more clearly as the dominant result, as evidenced by the errors being uniformly lower than those of the vanilla PINN. In addition, the learning curves suggest that the problem becomes less difficult with a substantially wider range of progress during training. Finally, for the subcritical cases, all models give similar solutions, except for $\alpha(t,r)$ produced by the vanilla PINN at $\phi_{0}=0.10625$. We also observe that every alternative model converges faster—roughly in half the training time—representing a clear advantage. However, in general, the main conclusion is that all methods perform reasonably well due to the relative “ease” of these problems: smooth gradients make them straightforward for any of the models to solve.
@@
It is important to note that these learning curves correspond to the training runs reported in table 1; thus, all methodological settings are the same, including the spatiotemporal sampling size $N_{\mathrm{t}}\times N_{\mathrm{r}}=128\times 128$. Several factors could be adjusted to improve results: increasing sampling density, fine-tuning $\epsilon_{\mathrm{t}}$ and $\epsilon_{\mathrm{r}}$, and / or refining the weights assigned to the initial and boundary conditions. For the critical case in particular, a more detailed analysis is needed, which we will present in the remainder of the paper.
@@
Robustness of ModPINN. Up to this point, we have reviewed a set of recent models in the literature, but it is worth putting more focus on our own proposal, which we have denoted as ModPINN. In essence, this architecture is no more than a modified PINN that replaces linear layers with quadratic ones, and whose main strength is an embedding introduced before the layer concatenation. As we have shown, this embedding leads to improved results for the supercritical case and, specifically, for the critical case $\phi_{0}=0.1125$, improving average performance and reducing variability. Taking into account that this model is the one giving the strongest results for gravitational collapse, it is necessary to carry out a more thorough examination of how various pre-training design choices influence its performance. Among these factors, we include the sampling size (that is, the number of spatiotemporal sampling points, $N_{\mathrm{t}}\times N_{\mathrm{r}}$) and the scale of the neural architecture, in particular the number of quadratic layers and the number of neurons per layer. Another parameter of interest that we examine here is the strength of the temporal causality, $\epsilon_{\mathrm{t}}$, while $\epsilon_{\mathrm{r}}$ will be constant and fixed at 10 for this analysis.
@@
  • The $10\times 75$ architecture—the deepest and widest configuration—yields the most stable overall results across all cases, much like the largest sampling size, $256\times 256$, which also exhibits the greatest stability. Consequently, the model clearly benefits from larger sampling grids and higher network capacities, providing robustness independent of the problem considered. In contrast, smaller networks display greater variability as $\epsilon_{\mathrm{t}}$ changes. This behavior is a clear indication of the inherent physical complexity of the problem.
  • In general, the motivation behind the enforcement of temporal causality is to prevent PINNs from “looking ahead” in time, that is, from resolving late-time points before early-time ones [Wang_2024]. This approach can be helpful in problems with a strong dependence on initial conditions, where the evolution of the system is fully determined by them. Consequently, it is reasonable to expect that in a smooth, easy-to-solve case like $\phi_{0}=0.1$, causality enforcement will have limited impact, which is essentially what we observe. For this problem, there is no clear trend in $l_{2}$ as $\epsilon_{\mathrm{t}}$ increases, because the solutions with $\epsilon_{\mathrm{t}}=0$ are generally good and any observed improvements are not substantial.
  • Following the same logic, for the critical case $\phi_{0}=0.1125$, we observe a stronger dependence on $\epsilon_{\mathrm{t}}$, although the errors are generally higher due to the steep highly localized gradients mentioned. We see that the model clearly benefits from enforcing causality, particularly for the $128\times 128$ case, where the error quickly drops for $\epsilon_{\mathrm{t}}=10^{2}$. Regarding the $\phi_{0}=0.125$ scenario—specifically with $256\times 256$ points and a $10\times 75$ network—the solution is markedly stable, showing only minor sensitivity to $\epsilon_{\mathrm{t}}$. By contrast, when smaller networks are used, the improvement becomes appreciable.
@@
Figure 4: Relative total $l_{2}$ error of the proposed ModPINN for the considered scalar-field initial-value problems (by column) as a function of the hyperparameter $\epsilon_{\mathrm{t}}$, which modulates the strength of the temporal causality. Each row corresponds to a different spatiotemporal grid resolution $(t,r)$, and each sub-figure reports results for three network sizes.
@@
In the two bottom panels, we depict how the $l^{2}$ errors vary with the hyperparameter $\Lambda$, which controls the “strength” with which spacetime points $(t,r)$ are moved towards regions exhibiting higher values of the normalized physical loss, as defined in (28). The lower-left panel reports this magnitude for the two subcritical cases ($\phi_{0}=0.1,0.10625$), while the lower-right panel presents the corresponding results for the critical and supercritical cases ($\phi_{0}=0.1125,0.125$). We observe that as $\Lambda$ increases within the range considered, the error shows a clear decreasing trend for the subcritical scenarios, with a particular reduction in variability at $\Lambda=10$ for the case $\phi_{0}=0.1$. For the critical scenario $\phi_{0}=0.1125$, the trend is also downward from $\Lambda=0$ to 1, with the error dropping from approximately $\sim 1.0$ to $\sim 0.6$. This constitutes a substantial improvement that should not be eclipsed by the use of a broader scale relative to the other panel. Regarding $\Lambda=10$ we observe a sudden increase in error for this case, indicating that the solutions deteriorate when this parameter is overly enlarged. This behavior is somewhat expected: the physics lies near the boundary of criticality, and the collocation points are required and used within a highly localized region of spacetime mainly determined by the black hole formation. Excessive and rapid relocation of the points can move some of them into undesirable regions, thereby devaluing their utility and, overall, leading to worse solutions. Finally, for the supercritical case with $\phi_{0}=0.125$, the trend is essentially flat. It is worth noting that the model under consideration here is ModPINN, which—consistent with table 1—provides the best performance for this regime with quite low variability. Consequently, this scenario is already well resolved without additional techniques, as the model alone effectively minimizes the loss. It is therefore unsurprising that the solution does not improve substantially when adaptive sampling is introduced. By contrast, performing the same analysis with models such as the vanilla PINN—where this case is less accurately captured—might reveal a decreasing trend under the said sampling methodology.
@@ 5 Conclusions
The work reported in this paper has revisited this challenging physical problem, focusing on solutions at and around the critical point. Our primary aim has been to demonstrate that neural network-based methodologies, such as Physics-Informed Neural Networks (PINNs), can serve as a viable and useful alternative to finite differences. To this end, we have considered four different initial-value problems for the field, using the amplitude of the initial kernel, $\phi_{0}$, as the control parameter, with the critical case included among them. We have addressed these cases using several recent methodologies from the literature and introduced our own, termed ModPINN, which has been tailored to this class of scenarios, including several novel aspects from recent works such as coordinate embedding, quadratic neural layers, and learnable activation functions. Our model delivers slightly better critical solutions with lower variability than other state-of-the-art approaches, even though all of them have led to very competitive predictions. Furthermore, by implementing a relatively simple algorithm that moves the spacetime sampling points $(t,r)$, we treat them as particles subject to a “force” with a prescribed acceleration modulated via the $\Lambda$ hyperparameter in (28), which indicates how quickly points migrate to regions with higher physical loss. The key consequence is that we are able to work with comparatively small sets, unlike traditional numerical methods that resolve black holes by brute force—highly densifying the mesh in spacetime regions that are known a priori to be most challenging.
@@
Irrespective of the specific models used, much of the substantive effort resides in the design of the physical loss. The boundary conditions for $r\gg 0$ introduced in (16), are formulated in terms of the derivatives of the physical variables, whose enforcement leads to the asymptotic behaviors that these variables are expected to have far from the source of perturbation. These conditions, as well as the mechanisms of temporal and spatial enforcement, have proved to be crucial and, in some sense, help to emulate the adaptive refinements of finite-difference schemes without incurring additional computational effort. In the analysis displayed in figure 4, we have examined the robustness of ModPINN with respect to several factors, including the density of sampling points and the width and depth of the network. We have found that wider and deeper architectures indeed facilitate better solutions, and they also perform relatively well on comparatively large spatiotemporal grids such as $256\times 256$ points—still small in comparison with those used in finite-difference schemes. Overall, these configurations exhibit greater robustness and smoother performance across independent trials. While further analysis could be considered, this allows us to highlight a notable advantage of PINNs: aside from hardware constraints, there is substantial flexibility to scale model size and sampling, and the resulting $N_{\mathrm{t}}\times N_{\mathrm{r}}$ grids in deep learning setups are orders of magnitude smaller than those typically required by finite differences.
@@
Across all models considered for the target problems, we can conclude that physics-informed deep-learning approaches are indeed capable of resolving gravitational collapse driven by a massless scalar field in a spherically symmetric, time-dependent spacetime. By judiciously selecting activation functions to restrict variables to their physically feasible ranges and by formulating boundary conditions in derivative form, we obtain solutions close to the ground truth. While there is still room for improvement, the results indicate that these models are progressing in the right direction. Nevertheless, a PINN can be seen as an alternative (or surrogate) to a finite-difference numerical solver. Broadly speaking, PINN-type models employ the governing PDEs as components of a constructed loss function—alongside additional information—to solve the physical problem at hand. Consequently, once the initial scalar field is specified, the full evolution of $\phi$, $\alpha$, and $C$ over $(t,r)$ is determined. Thus the problem becomes deterministic; gravitational collapse may or may not occur depending on that initial configuration. Therefore, this imposes certain limitations when investigating aspects such as the universality of the exponent $\gamma$ in the scaling law of the black hole mass $M_{\mathrm{BH}}$ near criticality, as well as the generality of the proportionality constant $c_{f}$ with respect to the family of the initial field. To examine these aspects—and to study the self-similarity of solutions as $p\to p^{*}$—we would need access to a broad set of initial condition sets and their full corresponding solutions. Therefore, addressing these questions will require a different point of view—one that allows us to traverse families of initial conditions with full flexibility and to generate data with relative ease. In this sense, the present project paves the way for future studies that can probe the aforementioned issues, extend to alternative spacetime metrics and matter models, and tackle more complex scenarios in which finite-difference methods could be either inapplicable or impractical due to prohibitive computational costs.
@@ 6 Acknowledgments
This work has been partially funded by the Valencian Government grant with Reference Number CIAICO/2024/111; the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call-QuantumSpain project, and the European Union through the Recovery, Transformation and Resilience Plan—NextGenerationEU within the framework of the Digital Spain 2025 Agenda. This work has been also supported by the Spanish Agencia Estatal de Investigación (grant PID2024-159689NB-C21) funded by MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe, by the Prometeo excellence programme grant CIPROM/2022/49 funded by the Generalitat Valenciana, and by the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA2021-SE-01 Grant No. NewFunFiCO-101086251. MWC is supported by NSERC. R. RdA is supported by PID2020-113644GB-I00 from the Spanish Ministerio de Ciencia e Innovación and by PROMETEO/2022/69 funded by the Generalitat Valenciana.